OpenStudy (anonymous):

If 6y+3x+8xy-94=0 and x+5=13, then y+4=?

OpenStudy (anonymous):

6y+24+64y-94=0 70y=70 y=1 Soo, Y+4=1+4=5

OpenStudy (anonymous):

121